# Stable electron-irradiated [1-13C]alanine radicals for metabolic imaging with Dynamic Nuclear Polarisation Short Title: Electron-irradiated radicals for DNP

### v2

@@ 6.2 DNP
DNP was performed either using an Oxford Instruments HyperSense polariser at 3.35 T or a customised SpinAligner 6.7 T polariser. In both cases either microwave sweep profiles were obtained by iterating over microwave irradiation frequency, or polarisation build-up curves in which a low flip angle pulse was repeatedly played and polarisation monitored. The carbon Larmor frequency at both field strengths were 35.96048861 MHz and 71.816528 MHz respectively. Amplitudes obtained and reported as signal during microwave sweeps or build-up curves are integrated over the well-defined NMR peak present in the 13C NMR signal as is commonly performed with DNP.
@@
$y=P\left(A\left(1-e^{-\frac{\text{time}}{T}}\right)+\left(1-A\right)\left(1-e^{-\frac{\text{time}}{\tau}}\right)\right)$ (S1)
where $P$ is the build-up constant, $A$ and 1A1-A are the relative weightings of the fast and slow components of the polarisation build-up, and $T$ and $\tau$ are the fast and slow components of the build-up times.
For the HyperSense, microwave frequency sweeps were acquired for each sample between 93.75 to 94.195 GHz in 10 MHz steps with a two minute build-up per frequency step. The microwave power was held constant at 100 mW for all microwave frequency sweeps and build-ups. The build-up of polarisation was subsequently monitored at the peak frequency of the microwave frequency sweep every three minutes using a low flip-angle readout. A 3.35 T frequency sweep is shown in Figure S2 together with the corresponding sweep for an alanine + OX063 sample. A small potential enhancement lobe may be seen at 93.970 GHz. Black dashed lines denote equivalent features in the 6.7 T microwave sweep curves, scaled exactly for field. Given the considerably less sophisticated microwave and RF architecture of the HyperSense compared to the SpinAligner, the possibility that this lobe at 93.970 GHz is artefactual remains, although it was not observed in the control sample.
Figure S2: The frequency sweep obtained at 3.35 T for both alanine and trityl radical and 100 kGy irradiated alanine. The parameters noted “LL” (lower limit), “UL” (upper limit) and “COM” (centre of mass) are presented for illustrative comparisons with higher field DNP data, and have been defined phenomenologically as the lower, centre and upper region of features in the 6.7 T curve scaled for the exact difference, i.e. the corresponding down-scaled values of [187.77, 187.95, 188.15][187.77,\,187.95,\,188.15] GHz at 3.358 T.
Based on previous optimisation work and the use of a narrow line width radical, the SpinAligner was used without microwave modulation for the non-irradiated samples. For the irradiated samples, microwave modulation was applied over a range of 25 MHz about the given centre frequency, sinusoidally varied with the frequency of microwave irradiation occurring at a rate of 1 kHz. This applies a weak modulation and attempts to invert a greater proportion of electron spins; it also is recommended by the microwave source manufacturer in order to promote greater thermal stability of their system.
@@
The NMR spectra recorded during microwave frequency sweeps and polarisation build-ups were quantified by peak integration in MATLAB (ver. 2023a; The Mathworks, Inc). Polarisation estimates were obtained in the liquid state both hyperpolarised and at thermal equilibrium, and then back-calculated using the measured $T_{1}$ and known time of dissolution to estimate solid-state polarisation; for samples with low nuclear polarisation in the solid state (where relative error post dissolution is likely higher or sufficiently high to be inaccurate) a direct linear scaling of ADC values obtained at constant gain with a known polarisation reference were used.
6.3 Power characterisation at 6.7 T
Figure S3: (A) Frequency sweep profiles comparing 10 mW and 20 mW microwave power showing preserved asymmetric enhancement, and marginally enhanced build-up. (B) Build-up profiles demonstrating increased absolute signal at higher power with largely unchanged kinetics. (C–G) Systematic investigation of frequency modulation parameters showing optimal conditions near 100 Hz modulation rate with preserved biexponential build-up characteristics across all conditions tested.
To address potential concerns regarding the optimisation of microwave irradiation conditions, we performed extensive calibration experiments investigating the effects of microwave power, and, for our source in the SpinAligner, frequency modulation and pulse timing parameters on the DNP enhancement profiles. These experiments were essential given the high-field (6.7 T, 188 GHz) operation of our system, where microwave field inhomogeneity within the sample cavity becomes increasingly problematic (?). This is because the SpinAligner’s microwave cavity (diameter 28 mm, height 30 mm) is highly overmoded at 188 GHz, precluding meaningful quality factor measurements that may come with unequal power deposition into the sample.
Additionally, owing both to the closed-cycle cryogenic nature of the machine and additional hardware constraints, continuous-wave irradiation was not feasible due to thermal instabilities in both the microwave source and the sample environment. This necessitated the pulsed modulation approach, which while effective for our current studies, should not be confused with (and, indeed, precludes) the investigation of coherent polarisation transfer mechanisms such as TOP-DNP (?) or NOVEL sequences (?, ?) that might further optimise the enhancement – our modulations are orders of magnitude slower than the electron Larmor frequency.
Systematic power-dependent measurements were performed on 70 kGy electron-irradiated 13C-alanine samples using our “standard conditions” that were used in the majority of this work (25 MHz amplitude modulation, 100 Hz frequency modulation, both detailed subsequently). Figure S3A shows frequency sweep profiles acquired at 10 mW and 20 mW output power, demonstrating that the characteristic asymmetric enhancement profile is preserved across this power range. The build-up kinetics (Figure S3B) show increased absolute signal intensity at higher power, consistent with more efficient microwave saturation of the electron spin transitions, but the fundamental biexponential character and relative time constants remain unchanged.
The SpinAligner system employs pulsed microwave irradiation with frequency modulation to accommodate the broad EPR linewidth of the stable alanine radicals while avoiding continuous-wave heating limitations of the 188 GHz source. The user-specified parameters are detailed below:
  • AM (Amplitude Modulation): The frequency sweep bandwidth in MHz. This determines the range of microwave frequencies scanned during each pulse (e.g., 25 MHz spans $\pm$12.5 MHz around the center frequency).
  • FM (Frequency Modulation): The base modulation rate in Hz. This parameter, combined with hardware duty cycle constraints, determines the pulse repetition timing.
The relationship between these user parameters and actual pulse timing is governed by the microwave synthesiser hardware implementation. For combined PM+FM operation, the pulse timing follows:
$t_{\text{pulse on}}$ $=t_{\text{off,displayed}}+\frac{10^{6}~\mu s/s}{f_{\text{FM}}}-40~\mu s$ (S2)
$t_{\text{pulse off}}$ $=t_{\text{on,displayed}}+40~\mu s$ (S3)
where $f_{\text{FM}}$ is the FM parameter in Hz, and the displayed on/off times are hardware-inverted values read from the synthesiser. The term $10^{6}/f_{\text{FM}}$ represents the base period in microseconds corresponding to the FM rate.
For our standard conditions (AM = 25 MHz, FM = 100 Hz), these equations yield:
$t_{\text{pulse on}}$ $=\frac{10^{6}}{100}-40=9960~\mu s\approx 9963~\mu s\text{ (measured)}$ (S4)
$t_{\text{pulse off}}$ $=48139~\mu s\text{ (measured)}$ (S5)
This corresponds to a duty cycle of approximately 17% and a total cycle time of  58 ms. The discrepancy between the theoretical base period (10 ms for 100 Hz) and actual cycle time reflects hardware-imposed timing constraints and thermal management requirements.
During each “on” period, the microwave frequency is swept linearly across the AM bandwidth (e.g. 25 MHz) in discrete steps. The number of frequency steps and dwell time per step are determined by the synthesiser’s internal timing, typically resulting in 25–50 discrete frequency points per sweep. This approach aims to ensure that the broad EPR linewidth ($\sim$15 mT FWHM, corresponding to $\sim$420 MHz at 6.7 T) is broadly saturated while maintaining sufficient power at each frequency point to drive polarisation transfer.
We systematically investigated the effects of both amplitude modulation bandwidth (25–100 MHz) and frequency modulation rate (100 Hz to 10 kHz) on the enhancement characteristics. Figure S3C–G summarises the parametric dependence of the biexponential build-up characteristics on modulation frequency. Although very preliminary values and potentially specific to our hardware, the build-up constant $P$ (panel C) shows optimal values near 100 Hz modulation, with diminishing returns at higher frequencies. This behavior likely reflects the interplay between the EPR linewidth ($\sim$15 mT FWHM), the frequency sweep range (25 MHz), and the electron spin relaxation times under our experimental conditions.
Importantly, the relative weightings of fast and slow build-up components (panels D and E) and their respective time constants (panels F and G) show some systematic but modest variations with modulation parameters. The fast component time constant $T$ remains in the range 300–400 s across all conditions tested, while the slow component $t$ varies between 0.5–2 hours (the upper end of the build-up experiment tested). We believe this rules out, for example, a finite microwave bandwidth leading to differential enhancement, and we further note that the pyruvic acid frequency sweep curve under these conditions is a well-resolved bimodal curve of equal intensity within this region.
6.4 Stability analysis
Samples of irradiated alanine and glycerol were created on or the day after irradiation, and stored in borosilicate glass haematocrit tubes, themselves stored in a falcon tube. These were stored in separate home-made desiccators consisting of monolayer of silica desiccant gel beads (Sigma Aldrich, UK) in a plastic ‘tupperware’ self-sealing container. Three repeated EPR measurements were obtained by inserting the haematocrit tubes inside an EPR tube, and were rotated by hand within the machine between acquisitions. EPR spectra were obtained at weeks 0, 2, 4, and 16 after irradiation. A simple linear regression approach was used for 30, 50 and 70 kGy irradiated samples dispersed in alanine; there was no significant change with time ($p$ = 0.14, 0.085, 0.18, respectively) and the goodness of fit was poor ($R^{2}=0.20$, $0.27$, and $0.17$ respectively). We interpret this as providing evidence of stability as there was no significant change of EPR signal intensity with time.
The raw data behind the stability plot is given below, in Table S1
Table S1: Spin-count data for EPR stability analysis of three replicate samples each of irradiated alanine dispersed in glycerol, measured over time. This is plotted graphically in Fig. 2B.
Week 30 kGy 50 kGy 70 kGy
0 3340 2910 3550 5590 4100 3370 6210 5420 4970
2 5070 2110 2800 5590 4740 4000 5070 4070 5460
4 2120 4030 4580 4990 6190 3740 5900 5390 4740
16 2780 2130 1890 4300 2360 3100 5100 3760 5040
@@ 6.5 Polarisation characterisation
This is potentially because of the competing effects of increased polarisation transfer efficiency and enhanced nuclear relaxation at higher radical concentrations. The shortened fast component likely reflects more efficient initial polarisation transfer due to increased electron-nuclear spin coupling, while the extended slow component may indicate limitations in long-range spin diffusion as paramagnetic centres become more densely packed. At higher radical concentrations, the shortened inter-radical distances could impede the cooperative network of nuclear spins necessary for efficient polarisation distribution throughout the sample, creating isolated ‘islands’ of polarisation around each radical centre. This spatial heterogeneity in the polarisation landscape would manifest as different characteristic time constants in the bi-exponential build-up curve.
@@
Table S2: Where $T$ is the fast component of the build-up time and $t$ is the slow component of the build-up time.
@@
6.6 Imaging experiment details
The EPSI sequence used is described in detail elsewhere (?) and was intended to both show the spatial localisation of the hyperpolarised contrast agent and additionally be sensitive (i.e. able to resolve) any labelled impurity peaks that may be present. The acquisition matrix size was 16 $\times$ 16 (spatial) $\times$ 256 (spectral); in an $\times80$ mm2 field of view; slice thickness: 30 mm; spectral bandwidth: 5000 Hz; for a spectral resolution of 19.5 Hz/point prior to reconstruction. Data were reconstructed with zerofilling by a factor of two in all dimensions. The TR was 16.0 seconds; TE: 2.642 ms; with a 15 flip angle. Eight images were acquired.
6.7 EPR Supplementary Methods
@@
Quantitative EPR spectra were recorded on an X-band continuous wave (CW) EPR spectrometer (EMXmicro, Bruker Biospin GMBH) using an ER 4122SHEQ resonator at room temperature; low temperature measurements, including electronic relaxation constants $T_{1e}$ and measured $T_{2e}$ (known as $T_{m}$) measurements were performed at 5 K on an X-band pulsed E680 spectrometer. Magnesium Oxide (MgO) powder, a common standard reference in EPR, (?) was used to account for any field offset in the experimental data, arising on physical grounds, which was calculated at $\sim 0.6$ mT. Furthermore, it is known that the signal amplitude is proportional to $\chi_{m}Q\sqrt{\text{Microwave power}}$ in EPR and these parameters were controlled appropriately. (?)
Samples that comprised dry powder were directly loaded into 4 mm thin wall quartz EPR tubes and their mass recorded. For repeated measurements, samples were rotated about their own axis by approximately 25 to 35  to ensure adequate powder averaging. Liquid state samples were first loaded in 2 mm diameter capillary tubes which in turn were loaded into the EPR tubes.
All solid state samples used the standard set of parameters listed below where only frequency changed between samples. The frequency has been normalised to for all data sets presented. Scans other than at 5 K were performed at room temperature. In brief, spectrometer parameters were: microwave attenuation = 36 dB; microwave power = 0.05024 mW; scans = 1; centre field = 333.8 mT; sweep width = 33 mT; sweep direction = up; sweep time = 40.04 s; time constant = 20.48 ms; conversion time = 14.46 ms; modulation amplitude = 1.2 G; modulation Frequency = 100 kHz; sweep rate = 0.816 mT/s. Echo Detected field sweeps for gamma irradiated alanine were recorded at 5 K using a standard Hahn echo pulse sequence ($\frac{\pi}{2}-\tau-\pi-\tau-\text{detect}$) with $\pi/2$ pulse length of 16 ns, $\pi$ pulse length of 32 ns and $\tau=300$ ns, a shot repetition rate of 1.024 s was used corresponding to a sweep rate of 0.2 mT/s with 1 shot per point and one average used. For ease of comparison to CW data this field swept data is presented in the derivative form in the main text. $T_{1e}$ was measured by fitting to a functional form $y=c-ae^{-\frac{x}{b}}$, where $b$ is the $T_{1}$ time, using the “fit” function in Matlab, performing a least-squares fit. The field sweep was for the sample containing OX063 was measured using free induction decay detection with a 1000 ns pulse followed by a detection period of 600 ns with a shot repetition rate of 0.2 s with 2 shots per point and a sweep width of 3 mT corresponding a sweep rate of 0.06 mT/s. Considering a g-value of 2.00319 it was found that this system had a field offset of 0.624 mT and spectra have been corrected accordingly.
@@
For longitudinal studies observing the stability of radical species, EPR spin counts were recorded on the same day as their irradiation and then two weeks, four weeks, and 16 weeks post irradiation. Dry powder samples were prepared directly in the EPR tubes and rotated about their own axis as above; the mass of empty and filled tubes were recorded and monitored over time and data are corrected for the mass of samples present. Solutions containing glycerol were prepared as stock solutions which were used to fill haematocrit tubes which themselves were inserted into the EPR tubes. Best practices for undertaking quantitative EPR spectroscopy were followed, such as the use of an appropriate resonator, and reproducible packing metrics with impurity controlled materials. (?, ?) EPR data was analysed in Matlab. Spin counts were determined by double numerical integration of the EPR spectra. First, the raw spectra were baseline-corrected using a first-order correction, followed by integration to obtain absorption spectra. Q-value normalisation was performed to that using a reference Q-value of 8000 (obtained with the MgO sample) to account for variations in cavity loading. A polynomial fit (order 15) was then applied to specific field regions (317.5–333 mT and 337–350 mT) for secondary baseline correction. The final double integration value was calculated from the fully corrected absorption spectrum, with verification performed by comparing the numerical derivative of the absorption spectrum against the original intensity data. Full width at half maximum (FWHM) values were also calculated for each spectrum to characterise line broadening effects related to radical concentration and interactions.
@@
The $T_{1e}$ for gamma irradiated alanine at 5 K recorded at the center of the observed signal was measured using an inversion recovery experiment with the sequence $\pi-t-\frac{\pi}{2}-\tau-\pi-\tau-\text{detect}$, with $\pi/2$ pulse length of 16 ns, $\pi$ pulse length of 32 ns and $\tau=$ 300 ns. The measured $T_{1}$ was greater than the allowed maximum shot repetition rate of the spectrometer and thus the command tsleep was used to generate the inter pulse delay t which gave an effective shot repetition rate of 60 s. The measured value of $T_{1e}$ for the SAR at 5 K in glycerol was 7.59 s, and $T_{m}\approx 332~ns$. For comparison, a “literature prep” (non-irradiated [1-13C]L-alanine, 18.94M NaOH, DMSO, 15 mM OX063 and 0.3 mM Dotarem gadolinium as Gd(III)) had $T_{1e}=39~ms$ (using a standard inversion recovery with a shot repetition rate of 0.25 s), Tm=378nsTm=378~ns. A “literature prep without gadolinium”, i.e. just OX063, using a shot repetition rate of 9 s and tsleep had $T_{1e}=1.41~s$ and $T_{m}=370~ns$. These fits are shown below, in Fig. S4. Raw data is provided as further SI to this manuscript.
Figure S4: Electronic $T_{1e}$ and $T_{m}$ measurement data and fits, obtained at 5 K, for A: irradiated 100 kGy alanine; B: a “literature prep” trityl radical without gadolinium; and C: a “literature prep with gadolinium” sample.
@@
6.8 Alanine SAR EPR Fitting and EPR Simulation Details
To try to fit the EPR data obtained of irradiated powder 13C alanine, a detailed simulation was undertaken in EasySpin using appropriate previously published hyperfine constants and $g$-tensors for the radical R1 only, which is responsible for the majority of the signal recorded via EPR. This was motivated by a desire to have a single radical species for further large-scale spin-density simulations (although we note that a semi-empirical approach was additionally undertaken, using purely the experimental EPR spectrum scaled for field). The initial fitting model included a spin system with an $S=1/2$ electron and a $g$-tensor consistent with the R1 alanine radical ($g=[2.0041,2.0034,2.0024]$). The model incorporated hyperfine couplings for five protons (four $\beta$-protons and one $\alpha$-proton) with starting values derived from NA alanine, plus the 13C nucleus . These are stated in full in the attached analysis code in Data S2. We employed the pepper function provided by EasySpin (?) with a hybrid method for orientational averaging, using an adaptive grid approach with an initial grid size of [10 2] and a minimum grid size of 2. This was progressively refined throughout the optimisation process. Isotropic line broadening from unresolved interactions was modelled with a peak-to-peak linewidth (lwpp) parameter of 0.45 mT.

### v3

@@ 6.2 DNP
DNP was performed either using an Oxford Instruments Hypersense polariser at 3.35 T or a customised SpinAligner 6.7 T polariser. In both cases either microwave sweep profiles were obtained by iterating over microwave irradiation frequency, or polarisation build-up curves in which a low flip angle pulse was repeatedly played and polarisation monitored.
@@
$y=P(A(1-e^{-\frac{\text{time}}{T}})+(1-A)(1-e^{-\frac{\text{time}}{t}}))$ (S1)
where $P$ is the build-up constant, $A$ and 1A1-A are the relative weightings of the fast and slow components of the polarization build-up, and $T$ and $t$ are the fast and slow components of the build-up times.
For the Hypersense, microwave frequency sweeps were acquired for each sample between 93.75 to 94.195 GHz in 5 MHz steps with a two minute build-up per frequency step. The microwave power was held constant at 100 mW for all microwave frequency sweeps and build-ups. The build-up of polarisation was subsequently monitored at the peak frequency of the microwave frequency sweep every three minutes using a low flip-angle readout.
Based on previous optimization work and the use of a narrow line width radical, the SpinAligner was used without microwave modulation for the non-irradiated samples. For the irradiated samples, microwave modulation was applied over a range of 25 MHz about the given centre frequency, sinusoidally varied with the frequency of microwave irradiation occurring at a rate of 1 kHz. This applies a weak modulation and attempts to invert a greater proportion of electron spins; it also is recommended by the microwave source manufacturer in order to promote greater thermal stability of their system.
@@
The NMR spectra recorded during microwave frequency sweeps and polarization build-ups were quantified by peak integration in MATLAB (ver. 2023a; The Mathworks, Inc). Polarization estimates were obtained in the liquid state both hyperpolarised and at thermal equilibrium, and then back-calculated using the measured $T_{1}$ and known time of dissolution to estimate solid-state polarisation; for samples with low nuclear polarization in the solid state (where relative error post dissolution is likely higher or sufficiently high to be inaccurate) a direct linear scaling of ADC values obtained at constant gain with a known polarisation reference were used.
@@ 6.3 Polarisation characterisation
This is potentially because of the competing effects of increased polarization transfer efficiency and enhanced nuclear relaxation at higher radical concentrations. The shortened fast component likely reflects more efficient initial polarization transfer due to increased electron-nuclear spin coupling, while the extended slow component may indicate limitations in long-range spin diffusion as paramagnetic centres become more densely packed. At higher radical concentrations, the shortened inter-radical distances could impede the cooperative network of nuclear spins necessary for efficient polarization distribution throughout the sample, creating isolated ‘islands’ of polarization around each radical centre. This spatial heterogeneity in the polarization landscape would manifest as different characteristic time constants in the bi-exponential build-up curve.
@@
Table S1: Where $T$ is the slow component of the build-up time and $t$ is the fast component of the build-up time.
@@
6.4 Spin count procedure and supplementary data
@@
Quantitative EPR spectra were recorded on an X-band CW EPR spectrometer (EMXmicro, Bruker Biospin GMBH) at room temperature; electronic $T_{2}$ measurements were performed at 5 K on an X-band E680 spectrometer. Magnesium Oxide (MgO) powder, a common standard reference in EPR, was used to account for changes in the magnetic field between experiments due to differences in sample positioning and resonator $Q$ values; it is known that the signal amplitude is proportional to $\chi_{m}Q\sqrt{\text{Microwave power}}$ in EPR and these parameters were controlled appropriately. (?) Samples that comprised dry powder were directly loaded into 4 mm thin wall quartz EPR tubes and their mass recorded. For repeated measurements, samples were rotated about their own axis by approximately 25 to 35  to ensure adequate powder averaging. Liquid state samples were first loaded in 2 mm diameter capillary tubes which in turn were loaded into the EPR tubes.
@@
For longitudinal studies observing the stability of radical species, EPR spin counts were recorded on the same day as their irradiation and then two weeks, four weeks, and 16 weeks post irradiation. Dry powder samples were prepared directly in the EPR tubes and rotated about their own axis as above; the mass of empty and filled tubes were recorded and monitored over time and data are corrected for the mass of samples present. Solutions containing glycerol were prepared as stock solutions which were used to fill haematocrit tubes which themselves were inserted into the EPR tubes. Best practices for undertaking quantitative EPR spectroscopy were followed, such as the use of an appropriate resonator, and reproducible packing metrics with impurity controlled materials. (?, ?) EPR data was analysed in Matlab. Spin counts were determined by double numerical integration of the EPR spectra. First, the raw spectra were baseline-corrected using a first-order correction, followed by integration to obtain absorption spectra. Q-value normalization was performed to that using a reference Q-value of 8000 (obtained with the MgO sample) to account for variations in cavity loading. A polynomial fit (order 15) was then applied to specific field regions (317.5–333 mT and 337–350 mT) for secondary baseline correction. The final double integration value was calculated from the fully corrected absorption spectrum, with verification performed by comparing the numerical derivative of the absorption spectrum against the original intensity data. Full width at half maximum (FWHM) values were also calculated for each spectrum to characterize line broadening effects related to radical concentration and interactions.
@@
For use with quantitative modelling studies, electronic $T_{1}$ and $T_{m}$s were collected via use of an inversion recovery pulse sequence ($T_{1}$) and $T_{m}$ (approximately $T_{2}$) by a 2-Pulse Echo Decay method. An approximate estimate of the electronic $T_{1}$ of the SAR at 5 K in glycerol was 12.4 s and $T_{m}\approx 371~ns$. For comparison, a “literature prep” (non-irradiated [1-13C]L-alanine, 18.94M NaOH, DMSO, 15mM OX063 and 0.3mM Dotarem gadolinium) had $T_{1e}=41~ms$, Tm=222nsTm=222~ns, “literature prep without gadolinium” had $T_{1e}=4.6~s$ and $T_{m}=181.6~ns$.
@@
6.5 Alanine SAR EPR Fitting
To try to fit the EPR data obtained of irradiated powder 13C alanine, a detailed simulation was undertaken in EasySpin using appropriate previously published hyperfine constants and $g$-tensors for the radical R1 only, which is responsible for the majority of the signal recorded via EPR. The initial fitting model included a spin system with an $S=1/2$ electron and a $g$-tensor consistent with the R1 alanine radical ($g=[2.0041,2.0034,2.0024]$). The model incorporated hyperfine couplings for five protons (four $\beta$-protons and one $\alpha$-proton) with starting values derived from NA alanine, plus the 13C nucleus (summarised below, in Code Listing 1). We employed the pepper function provided by EasySpin (?) with a hybrid method for orientational averaging, using an adaptive grid approach with an initial grid size of [10 2] and a minimum grid size of 2. This was progressively refined throughout the optimisation process. Isotropic line broadening from unresolved interactions was modelled with a peak-to-peak linewidth (lwpp) parameter of 0.45 mT.